# Quasi-linear equation $\Delta_{p}v+av^{q}=0$ on manifolds with integral bounded Ricci curvature and geometric applications

### v2

@@ 2. Volume estimate
$\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}\mathcal{L}(f^{\alpha})\geq&\frac{(2\alpha-1)b^{2}e^{2cu}f^{2-p}}{(2\alpha-1)(n-1)+p-1}-(n-1){\rm Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\\ &+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{f^{2}}{n-1}.\end{split}$ (3.20)
@@
$\begin{split}&\frac{(2\alpha-1)b^{2}e^{2cu}f^{2-p}}{(2\alpha-1)(n-1)+p-1}+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}\\ \geq&-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}f^{2}.\end{split}$ (3.21)
@@
$\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}\mathcal{L}(f^{\alpha})\geq$ $\left(\frac{1}{n-1}-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}\right)f^{2}$
$-{\rm Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|.$
@@
$\mathcal{L}(f^{\alpha})\geq$ $2\beta_{n,p,q,\alpha}\alpha f^{\alpha+\frac{p}{2}}-2\alpha{\rm Ric}_{-}f^{\alpha+\frac{p}{2}-1}-a_{1}\alpha f^{\alpha+\frac{p}{2}-\frac{3}{2}}|\nabla f|,$
@@
3.3. Approximation procedure and key integral inequality
@@
$t\in\left(\frac{a_{1}^{2}}{a_{2}\beta_{n,p,q,\alpha}},\,\,\infty\right),$ (3.22)
@@
$\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}+\gamma(\alpha,p,t)\int_{\Omega}f^{\theta}|\nabla\eta|^{2},$ (3.23)
@@
$-\int_{\Omega}\bm{\langle}f^{p/2-1}\nabla f^{\alpha}+(p-2)f^{p/2-2}\bm{\langle}\nabla f^{\alpha},\nabla u\bm{\rangle}\nabla u,\,\nabla\psi\bm{\rangle}$
$\geq$ $2\beta_{n,p,q,\alpha}\alpha\int_{\Omega}f^{\alpha+\frac{p}{2}}f^{t}_{\epsilon}\eta^{2}-2\alpha\int_{\Omega}{\rm Ric}_{-}f^{\alpha+\frac{p}{2}-1}f^{t}_{\epsilon}\eta^{2}-a_{1}\alpha\int_{\Omega}f^{\alpha+\frac{p-3}{2}}f^{t}_{\epsilon}|\nabla f|\eta^{2}.$
@@
$\begin{split}&-\int_{\Omega}(\alpha tf^{\alpha+\frac{p}{2}-2}f^{t-1}_{\epsilon}|\nabla f|^{2}\eta^{2}+t\alpha(p-2)f^{\alpha+\frac{p}{2}-3}f^{t-1}_{\epsilon}\bm{\langle}\nabla f,\nabla u\bm{\rangle}^{2}\eta^{2})\\ &-\int_{\Omega}(2\eta\alpha f^{\alpha+\frac{p}{2}-2}f^{t}_{\epsilon}\bm{\langle}\nabla f,\nabla\eta\bm{\rangle}+2\alpha\eta(p-2)f^{\alpha+\frac{p}{2}-3}f^{t}_{\epsilon}\bm{\langle}\nabla f,\nabla u\bm{\rangle}\bm{\langle}\nabla u,\nabla\eta\bm{\rangle})\\ \geq&2\beta_{n,p,q,\alpha}\alpha\int_{\Omega}f^{\alpha+\frac{p}{2}}f^{t}_{\epsilon}\eta^{2}-2\alpha\int_{\Omega}{\rm Ric}_{-}f^{\alpha+\frac{p}{2}-1}f^{t}_{\epsilon}\eta^{2}-a_{1}\alpha\int_{\Omega}f^{\alpha+\frac{p-3}{2}}f^{t}_{\epsilon}|\nabla f|\eta^{2}.\end{split}$ (3.24)
@@
$f^{t-1}_{\epsilon}|\nabla f|^{2}+(p-2)f^{t-1}_{\epsilon}f^{-1}\bm{\langle}\nabla f,\nabla u\bm{\rangle}^{2}\geq a_{2}f^{t-1}_{\epsilon}|\nabla f|^{2},$ (3.25)
@@
$f^{t}_{\epsilon}\bm{\langle}\nabla f,\nabla\eta\bm{\rangle}+(p-2)f^{t}_{\epsilon}f^{-1}\bm{\langle}\nabla f,\nabla u\bm{\rangle}\bm{\langle}\nabla u,\nabla\eta\bm{\rangle}\geq-(p+1)f^{t}_{\epsilon}|\nabla f||\nabla\eta|.$ (3.26)
@@
$\begin{split}&2\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+a_{2}t\int_{\Omega}f^{\theta-2}|\nabla f|^{2}\eta^{2}\\ \leq&2\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}+a_{1}\int_{\Omega}f^{\theta-\frac{1}{2}}|\nabla f|\eta^{2}+2(p+1)\int_{\Omega}f^{\theta-1}|\nabla f||\nabla\eta|\eta.\end{split}$ (3.27)
Since $u\in W^{2,2}_{loc}(\Omega\setminus\Omega_{cr})\cap C^{1,\beta}(\Omega)$ and the measure of critical set $\Omega_{cr}$ is zero by a very recent result [2, Corollary 1.6], we have $f\in C^{\beta}(\Omega)$ and $|\nabla f|\in L^{2}_{loc}$, and hence the integrals in the above make sense.
@@
$a_{1}f^{\theta-\frac{1}{2}}|\nabla f|\eta^{2}\leq\frac{a_{2}t}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+\frac{a_{1}^{2}}{a_{2}t}f^{\theta+1}\eta^{2},$ (3.28)
@@
$2(p+1)f^{\theta-1}|\nabla f||\nabla\eta|\eta\leq\frac{a_{2}t}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+\frac{4(p+1)^{2}}{a_{2}t}f^{\theta}|\nabla\eta|^{2}.$ (3.29)
@@
$\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2}\int_{\Omega}f^{\theta-2}|\nabla f|^{2}\eta^{2}\leq 2\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}+\frac{4(p+1)^{2}}{a_{2}t}\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.$ (3.30)
@@
$\begin{split}\frac{1}{2}\left|\nabla\left(f^{\frac{\theta}{2}}\eta\right)\right|^{2}\leq&\left|\nabla f^{\frac{\theta}{2}}\right|^{2}\eta^{2}+f^{\theta}|\nabla\eta|^{2}\\ =&\frac{\theta^{2}}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+f^{\theta}|\nabla\eta|^{2}.\end{split}$ (3.31)
@@
$\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{2a_{2}t}{\theta^{2}}\int_{\Omega}\left|\nabla\left(f^{\frac{\theta}{2}}\eta\right)\right|^{2}$ (3.32)
$\leq$ $2\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}+\left(\frac{4(p+1)^{2}}{a_{2}t}+\frac{2a_{2}t}{\theta^{2}}\right)\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.$
@@
$\begin{split}&\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\\ \leq&\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}+\left(\frac{2(p+1)^{2}}{a_{2}t}+\frac{a_{2}t}{\theta^{2}}\right)\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.\end{split}$ (3.33)
@@
3.4. Local $L^{\theta\chi}$ bound of the gradient.
@@
$\beta(\alpha,p,t)\triangleq\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|{\rm Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}>0.$
@@
$\|f\|_{L^{\theta\chi}(B_{3R/4}(o))}\leq a_{3}\left(\frac{V}{R^{2(\theta+1)}}\right)^{\frac{1}{\theta}},$ (3.34)
@@
$\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}\leq\left\|{\rm Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}.$
@@
$\frac{\beta_{n,p,q,\alpha}}{2}\int_{B(o,R)}f^{\theta+1}\eta^{2}+\left(\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|{\rm Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\right)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}$ (3.35)
@@
$\begin{split}\gamma(\alpha,p,t)\int_{B(o,R)}f^{\theta}|\nabla\eta|^{2}\leq&\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{R^{2}}\int_{B(o,R)}f^{\theta}\eta^{\frac{2\theta}{\theta+1}}\\ \leq&\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{R^{2}}\left(\int_{B(o,R)}f^{\theta+1}\eta^{2}\right)^{\frac{\theta}{\theta+1}}V^{\frac{1}{\theta+1}}\\ \leq&\frac{\beta_{n,p,q,\alpha}}{2}\left[\int_{B(o,R)}f^{\theta+1}\eta^{2}+\left(\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{\beta_{n,p,q,\alpha}R^{2}}\right)^{\theta+1}V\right].\end{split}$ (3.36)
@@
$\left\|f^{\theta}\right\|_{L^{\chi}(B_{3R/4}(o))}\leq\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}(B_{R}(o))}.$ (3.37)
@@
$\left(\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|{\rm Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\right)>0,$
@@
$\left\|f^{\theta}\right\|_{L^{\chi}(B_{3R/4}(o))}\leq\frac{\beta_{n,p,q,\alpha}}{\beta(\alpha,p,t)}\left(\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{\beta_{n,p,q,\alpha}R^{2}}\right)^{\theta+1}V.$ (3.38)
@@
$2\left(\alpha+t+\frac{p}{2}\right)>\beta^{*}$
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}<C(n,p,q,\beta^{*})\mathbb{S}_{\chi}(M),$
@@
3.5. Local gradient estimate when ${\rm Ric}_{-}\in L^{\gamma}$ for some $\gamma>\chi/(\chi-1)$.
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Denote $\Lambda=\|{\rm Ric}_{-}\|_{L^{\gamma}(B_{1})}$ for some $\gamma>\chi/(\chi-1)$. Then, for any $r\leq 1$ when
@@
$\frac{\beta_{n,p,q,\alpha}}{2}\int_{B_{r}}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\int_{B_{r}}{\rm Ric}_{-}f^{\theta}\eta^{2}+\gamma(\alpha,p,t)\int_{B_{r}}f^{\theta}|\nabla\eta|^{2}.$ (3.39)
@@
$\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}\leq\Lambda\|f^{\frac{\theta}{2}}\eta\|_{L^{\frac{2\gamma}{\gamma-1}}}^{2}.$ (3.40)
@@
$\int_{\Omega}{\rm Ric}_{-}f^{\theta}\eta^{2}\leq 2\Lambda\varepsilon^{2}\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2\chi}}^{2}+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2}}^{2}.$
@@
$\frac{\beta_{n,p,q,\alpha}}{2}\int_{B_{r}}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}$ (3.41)
@@
$\|f\|_{L^{\theta\chi}(B_{3r/4}(o))}\leq a_{3}\left(\frac{V}{r^{2(\theta+1)}}\right)^{\frac{1}{\theta}},$ (3.42)
@@
$\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\left(\gamma(\alpha,p,t)+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\right)|\nabla\eta|_{L^{\infty}}^{2}\int_{\Omega}f^{\theta}.$ (3.43)
@@
$\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\zeta(t)|\nabla\eta|_{L^{\infty}}^{2}\int_{\Omega}f^{\theta}.$ (3.44)

### v3

@@ 2. Volume estimate
(3.20) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq&\frac{(2\alpha-1)b^{2}e^{2cu}f^{2-p}}{(2\alpha-1)(n-1)+p-1}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\\ &+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{f^{2}}{n-1}.\end{split}$
@@
(3.21) $\begin{split}&\frac{(2\alpha-1)b^{2}e^{2cu}f^{2-p}}{(2\alpha-1)(n-1)+p-1}+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}\\ \geq&-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}f^{2}.\end{split}$
@@
$\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq$ $\left(\frac{1}{n-1}-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}\right)f^{2}$
$-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|.$
@@
${\mathscr{L}}(f^{\alpha})\geq$ $2\beta_{n,p,q,\alpha}\alpha f^{\alpha+\frac{p}{2}}-2\alpha\mathrm{Ric}_{-}f^{\alpha+\frac{p}{2}-1}-a_{1}\alpha f^{\alpha+\frac{p}{2}-\frac{3}{2}}|\nabla f|,$
@@
3.C. Approximation procedure and key integral inequality
@@
(3.22) $t\in\left(\frac{a_{1}^{2}}{a_{2}\beta_{n,p,q,\alpha}},\,\,\infty\right),$
@@
(3.23) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\gamma(\alpha,p,t)\int_{\Omega}f^{\theta}|\nabla\eta|^{2},$
@@
$-\int_{\Omega}\langle f^{p/2-1}\nabla f^{\alpha}+(p-2)f^{p/2-2}\langle\nabla f^{\alpha},\nabla u\rangle\nabla u,\,\nabla\psi\rangle$
$\geq$ $2\beta_{n,p,q,\alpha}\alpha\int_{\Omega}f^{\alpha+\frac{p}{2}}f^{t}_{\epsilon}\eta^{2}-2\alpha\int_{\Omega}\mathrm{Ric}_{-}f^{\alpha+\frac{p}{2}-1}f^{t}_{\epsilon}\eta^{2}-a_{1}\alpha\int_{\Omega}f^{\alpha+\frac{p-3}{2}}f^{t}_{\epsilon}|\nabla f|\eta^{2}.$
@@
(3.24) $\begin{split}&-\int_{\Omega}(\alpha tf^{\alpha+\frac{p}{2}-2}f^{t-1}_{\epsilon}|\nabla f|^{2}\eta^{2}+t\alpha(p-2)f^{\alpha+\frac{p}{2}-3}f^{t-1}_{\epsilon}\langle\nabla f,\nabla u\rangle^{2}\eta^{2})\\ &-\int_{\Omega}(2\eta\alpha f^{\alpha+\frac{p}{2}-2}f^{t}_{\epsilon}\langle\nabla f,\nabla\eta\rangle+2\alpha\eta(p-2)f^{\alpha+\frac{p}{2}-3}f^{t}_{\epsilon}\langle\nabla f,\nabla u\rangle\langle\nabla u,\nabla\eta\rangle)\\ \geq&2\beta_{n,p,q,\alpha}\alpha\int_{\Omega}f^{\alpha+\frac{p}{2}}f^{t}_{\epsilon}\eta^{2}-2\alpha\int_{\Omega}\mathrm{Ric}_{-}f^{\alpha+\frac{p}{2}-1}f^{t}_{\epsilon}\eta^{2}-a_{1}\alpha\int_{\Omega}f^{\alpha+\frac{p-3}{2}}f^{t}_{\epsilon}|\nabla f|\eta^{2}.\end{split}$
@@
(3.25) $f^{t-1}_{\epsilon}|\nabla f|^{2}+(p-2)f^{t-1}_{\epsilon}f^{-1}\langle\nabla f,\nabla u\rangle^{2}\geq a_{2}f^{t-1}_{\epsilon}|\nabla f|^{2},$
@@
(3.26) $f^{t}_{\epsilon}\langle\nabla f,\nabla\eta\rangle+(p-2)f^{t}_{\epsilon}f^{-1}\langle\nabla f,\nabla u\rangle\langle\nabla u,\nabla\eta\rangle\geq-(p+1)f^{t}_{\epsilon}|\nabla f||\nabla\eta|.$
@@
(3.27) $\begin{split}&2\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+a_{2}t\int_{\Omega}f^{\theta-2}|\nabla f|^{2}\eta^{2}\\ \leq&2\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+a_{1}\int_{\Omega}f^{\theta-\frac{1}{2}}|\nabla f|\eta^{2}+2(p+1)\int_{\Omega}f^{\theta-1}|\nabla f||\nabla\eta|\eta.\end{split}$
Since $u\in W^{2,2}_{loc}(\Omega\setminus\Omega_{cr})\cap C^{1,\beta}(\Omega)$ and the measure of critical set $\Omega_{cr}$ is zero by a very recent result [3, Corollary 1.6], we have $f\in C^{\beta}(\Omega)$ and $|\nabla f|\in L^{2}_{loc}$, and hence the integrals in the above make sense.
@@
(3.28) $a_{1}f^{\theta-\frac{1}{2}}|\nabla f|\eta^{2}\leq\frac{a_{2}t}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+\frac{a_{1}^{2}}{a_{2}t}f^{\theta+1}\eta^{2},$
@@
(3.29) $2(p+1)f^{\theta-1}|\nabla f||\nabla\eta|\eta\leq\frac{a_{2}t}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+\frac{4(p+1)^{2}}{a_{2}t}f^{\theta}|\nabla\eta|^{2}.$
@@
(3.30) $\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2}\int_{\Omega}f^{\theta-2}|\nabla f|^{2}\eta^{2}\leq 2\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\frac{4(p+1)^{2}}{a_{2}t}\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.$
@@
(3.31) $\begin{split}\frac{1}{2}\left|\nabla\left(f^{\frac{\theta}{2}}\eta\right)\right|^{2}\leq&\left|\nabla f^{\frac{\theta}{2}}\right|^{2}\eta^{2}+f^{\theta}|\nabla\eta|^{2}\\ =&\frac{\theta^{2}}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+f^{\theta}|\nabla\eta|^{2}.\end{split}$
@@
(3.32) $\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{2a_{2}t}{\theta^{2}}\int_{\Omega}\left|\nabla\left(f^{\frac{\theta}{2}}\eta\right)\right|^{2}$
$\leq$ $2\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\left(\frac{4(p+1)^{2}}{a_{2}t}+\frac{2a_{2}t}{\theta^{2}}\right)\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.$
@@
(3.33) $\begin{split}&\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\\ \leq&\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\left(\frac{2(p+1)^{2}}{a_{2}t}+\frac{a_{2}t}{\theta^{2}}\right)\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.\end{split}$
@@
3.D. Local $L^{\theta\chi}$ bound of the gradient.
@@
$\beta(\alpha,p,t)\triangleq\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}>0.$
@@
(3.34) $\|f\|_{L^{\theta\chi}(B_{3R/4}(o))}\leq a_{3}\left(\frac{V}{R^{2(\theta+1)}}\right)^{\frac{1}{\theta}},$
@@
$\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}\leq\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}.$
@@
(3.35) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{B(o,R)}f^{\theta+1}\eta^{2}+\left(\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\right)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}$
@@
(3.36) $\begin{split}\gamma(\alpha,p,t)\int_{B(o,R)}f^{\theta}|\nabla\eta|^{2}\leq&\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{R^{2}}\int_{B(o,R)}f^{\theta}\eta^{\frac{2\theta}{\theta+1}}\\ \leq&\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{R^{2}}\left(\int_{B(o,R)}f^{\theta+1}\eta^{2}\right)^{\frac{\theta}{\theta+1}}V^{\frac{1}{\theta+1}}\\ \leq&\frac{\beta_{n,p,q,\alpha}}{2}\left[\int_{B(o,R)}f^{\theta+1}\eta^{2}+\left(\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{\beta_{n,p,q,\alpha}R^{2}}\right)^{\theta+1}V\right].\end{split}$
@@
(3.37) $\left\|f^{\theta}\right\|_{L^{\chi}(B_{3R/4}(o))}\leq\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}(B_{R}(o))}.$
@@
$\left(\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\right)>0,$
@@
(3.38) $\left\|f^{\theta}\right\|_{L^{\chi}(B_{3R/4}(o))}\leq\frac{\beta_{n,p,q,\alpha}}{\beta(\alpha,p,t)}\left(\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{\beta_{n,p,q,\alpha}R^{2}}\right)^{\theta+1}V.$
@@
$2\left(\alpha+t+\frac{p}{2}\right)>\frac{2\chi}{\chi-1}\quad\mbox{as}\,\,\chi\in\left(1,\,\frac{n}{n-2}\right)$
or
$2\left(\alpha+t+\frac{p}{2}\right)>\beta^{*}\geq n\quad\mbox{as}\,\,\chi=\frac{n}{n-2},$
@@
$\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}<C(n,p,q,\beta^{*})\mathbb{S}_{\chi}(M),$
@@
3.E. Local gradient estimate when $\mathrm{Ric}_{-}\in L^{\gamma}$ for some $\gamma>\chi/(\chi-1)$.
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Denote $\Lambda=\|\mathrm{Ric}_{-}\|_{L^{\gamma}(B_{1})}$ for some $\gamma>\chi/(\chi-1)$. Then, for any $r\leq 1$ when
@@
(3.39) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{B_{r}}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\int_{B_{r}}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\gamma(\alpha,p,t)\int_{B_{r}}f^{\theta}|\nabla\eta|^{2}.$
@@
(3.40) $\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}\leq\Lambda\|f^{\frac{\theta}{2}}\eta\|_{L^{\frac{2\gamma}{\gamma-1}}}^{2}.$
@@
$\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}\leq 2\Lambda\varepsilon^{2}\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2\chi}}^{2}+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2}}^{2}.$
@@
(3.41) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{B_{r}}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}$
@@
(3.42) $\|f\|_{L^{\theta\chi}(B_{3r/4}(o))}\leq a_{3}\left(\frac{V}{r^{2(\theta+1)}}\right)^{\frac{1}{\theta}},$
@@
(3.43) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\left(\gamma(\alpha,p,t)+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\right)|\nabla\eta|_{L^{\infty}}^{2}\int_{\Omega}f^{\theta}.$
@@
(3.44) $\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\zeta(t)|\nabla\eta|_{L^{\infty}}^{2}\int_{\Omega}f^{\theta}.$